# Factorisation conditions and causality for local measurements in QFT

### v1

@@ III.1.1 Eliminating the ‘impossible’
which is a coupling between the field and one probe, localised to some region $R_{i}$. Approaches to measurement theory in QFT in which the measurement process is modelled as a local scattering process and operations are induced by probes, include the detector-based measurement theory [Polo_Gomez_2022] and the Fewster-Verch (FV) framework in algebraic QFT [fewster2020quantum]. In the detector-based approach, one works with (a general class) of Hamiltonian densities that model the detector-field coupling e.g. mimicking the light-matter interaction, and the local $S$-matrix can be defined as the time-ordered exponential of these interaction Hamiltonians (typically the predictions of the models rely on perturbation theory, using the Dyson expansion of the $S$-matrix). In the algebraic approach, the formulation of the FV framework relies on the structure of abstract *-algebras. In this setting, the interaction between system and probe field is formulated as a morphism $\Theta$ of the uncoupled theory, and it is not necessarily implemented through the adjoint action of a unitary scattering map $S$, namely as $\Theta(\cdot)=S^{\dagger}(\cdot)S$. Note that the approach we are taking here, is more abstract than the first (since it does not rely on particular interaction Hamiltonians and the Dyson expansion), but more concrete than the latter, since we assume the existence of these local scattering maps that implement the scatteting morphism in a Hilbert space.
@@
Partial or exact resolution to the impossible measurements scenario have been offered respectively in the detector-based approach [benincasa2014quantum, deramon2021relativistic] and in the FV-framework [PhysRevD.103.025017]. In either case, an important ingredient of the resolution was an analogue causal factorisation property (32). An abstract version of causal factorisation is assumed to hold in the FV framework  [^10] The abstract version of the causal factorisation condition is that $\Theta_{\textsc{a+b}}=\Theta_{\textsc{b}}\circ\Theta_{\textsc{a}}$, which has to be proven for concrete models in this framework., while in [deramon2021relativistic] it was argued that that this property holds for a general class of Unruh-DeWitt-type detector models coupled locally to a coupled field over disjoint, causally orderable regions. Finally, all the measurement schemes considered in QFT correspond to so-called pre-measurement processes, involving only the unitary interaction between the field and the probe. Strictly speaking, the measurement is completed only once an outcome has been obtained, which would require, in turn, a measurement of the probe and so on, leading to an infinite regress, see [mandrysch2024quantumfieldmeasurementsfewsterverch]. In this work we ignore this final measurement stage and, following common jargon, refer to the pre-measurement stage itself as the “measurement process”.
@@
We should emphasise that in both approaches causal factorisation is assumed (or proven on a case-by-case basis) but, crucially, it is not sufficient for eliminating impossible measurements. More assumptions about the probe-field interactions in the coupling regions have to be made, i.e. the locality of the detector’s local current in the detector model approach [deramon2021relativistic]  [^11] In [deramon2021relativistic] the authors show that (seemingly local) Hamiltonian densities of the form $h(\mathsf{x})=\Lambda(\mathsf{x})J(\mathsf{x})\otimes\phi(\mathsf{x})$, where $J(\mathsf{x})$ is the current associated to a probe, can give rise to impossible measurements if the probe’s dynamical current is not microcausal (i.e. if $J(\mathsf{x}),J(\mathsf{y})]\neq 0$ for $\mathsf{x},\mathsf{y}$ is not spacelike separated) by using the Dyson expansion of $S=\mathcal{T}\text{exp}i\int h(\mathsf{x})\mathrm{d}\mathsf{x}$. , or the time-slice property of the coupled theory in the FV framework  [^12] The exact relation between continuous additivity and other dynamical axioms e.g. the time-slice property in the formal setting of (p)AQFT goes way beyond our purposes (for related results see e.g. [ruep_thesis_2022]). [PhysRevD.103.025017]. Our claim is that Eq (31) captures the properties of a causal probe-field coupling that arises from the additional assumptions needed in the concrete and abstract frameworks, while making it clear that this can be directly related to internal factorisation properties of the interactions. Specifically, we will see that the ability to appropriately factorise the interaction region into causally ordered subregions is sufficient for ruling out impossible, i.e. signalling, operations. In the next section, we will see that it is not necessary, since a weaker factorisation property can be used.
@@ III.2 Hierarchy of factorisation conditions
We have introduced two notions of factorisations of $S$-matrices, causal factorisation and continuous additivity, that respectively correspond to composing disjoint interactions or decomposing a single interaction. Overall, there is a hierarchy of spacetime factorisation properties, as explained in [Brunetti_2000]. In this section, we will briefly outline them here for completeness, and generalise our proof of no impossible operations to a weaker condition.
@@
One such weaker condition, called the Hammerstein property, is typically assumed in causal perturbation theory literature [rejzner2019localitycausalityperturbativeaqft], where it is taken as continuous additivity Eq (31) may be too strong for very singular interactions [Brunetti_2000]. The Hammerstein property is that
@@
Importantly, the Hammerstein property is also sufficient to rule out Sorkin problems, as we explicitly show in detail in the appendix VII.1 and sketch here.
@@
Looking at Figure 5, we can see a similar scenario to the proof for the stronger factorisation condition. As there, the supports of $f_{\pm}$ are spacelike to C,A respectively, and now $f_{0}$ is spacelike to A and C. This implies that $S[f_{+}+f_{0}]S[f_{0}]^{\dagger}$ commutes with $C$ and $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is localised spacelike to A. Following a very similar derivation to Eq (III.1.1) demonstrates that there is no signalling in a Sorkin scenario where $\Phi_{\textsc{b}}$ is induced by an $S$-matrix satisfying the Hammerstein property  [^13] Resolutions of the Sorkin problem in the abstract algebraic setting of the FV framework rely on the time-slice property of the coupled theory [PhysRevD.103.025017, papageorgiou2024eliminating]. At the level of the local $S$-matrix it has been shown that the time-slice property can be derived only from (36) and normalisation $S(0)=\mathsf{1}$ [chilian2009time], which explains why we were able to demonstrate the link between (36) and no impossible operations in appendix VII.1.. For the full details, see appendix VII.1.
@@ IV Operations induced by probes
In quantum mechanics, every unital, i.e. maps the identity to itself, CP map (UCP map) – that is, every non-selective operation on a quantum system – can be understood as arising from a unitary interaction between the system and an ancillary system, or probe, followed by discarding (tracing out) the probe. This is a consequence of the Stinespring dilation theorem, which ensures that any UCP transformation of the system can be “purified” by embedding it into a larger closed system that evolves unitarily. In the same vein, we will consider in this section the realisation of non-selective operations in QFT through interactions between the system (i.e., the field) and a probe. Then, a way to argue that unitary “kicks” which violate no-signalling are not physically realisable is to show that they cannot be implemented by any valid local interaction between the system and a probe (see also [muchverch]).
@@ IV.1 Induced Kraus operators in QFT
The attempts of modelling local operations in QFT required a departure from the asymptotic scattering paradigm. In a series of papers [hellwigkraus1969, hellwigkraus1970b] Hellwig and Kraus model relativistic quantum measurement as a local scattering process. Their motivation for introducing the so-called Kraus operators, as induced through the dynamical coupling to some probe, was to explicate the operational interpretation of the local algebras that was implicit in the axiomatic approaches to QFT [haag_kastler_1964]. Their measurement scheme can be viewed as predecessor of recently proposed measurement frameworks in QFT [Polo_Gomez_2022, fewster2020quantum] that involve some (abstract or concrete) representation of dynamics between system and probe (for historical context see [Fraser2023]).
@@
Kraus operators in scattering-based frameworks for local measurement in QFT have been derived previously in [SmithAlexander2017, Polo_Gomez_2022, PhysRevD.109.065024] for concrete Unruh-DeWitt-type models, and more recently in [mandrysch2024quantumfieldmeasurementsfewsterverch] for a concrete representation of the FV framework in a Hilbert space. Nevertheless, the operational consequences of the factorisation conditions at the level of Kraus operators have not been investigated. Here we investigate how these properties are captured by valid local scattering matrices and their (de)composition properties. Further, for certain concrete model for measurement of field amplitudes, we link these properties to a no-signalling property for the resulting Kraus operators.
@@ IV.2 A concrete model for local field measurements
We will use a simple model for field measurements (previously considered in [PhysRevD.109.065024]) where a scalar quantum field $\phi$ in Minkowski spacetime is coupled to a free pointer via $P$, with $[Q,P]=i$. We specify the following interaction
@@
Since we can pick any free Hamiltonian for the pointer that commutes with the interaction without much computational difficulty, we could in principle pick that of a relativistic scalar particle. However, to avoid any conceptual problems arising from interpreting the pointer as a relativistic particle—given the well known problems of causality and localisation in relativistic QM (e.g. Hegerfeldt’s [hegerfeldt2001particle] and Malament’s [malament1996defense] theorems)—we elect simply to set $H_{\text{pointer}}=0$, viewing the pointer as an abstract, degree of freedom, not tied to spacetime. This has no effect on the field Kraus operators, and the change to the $S$-matrix can be found easily. To emphasise this conceptual choice, we will continue to denote spacetime points by $\mathsf{x}$ etc, while the internal values of the pointer as $q$. The interaction in (44) leads to the following scattering matrix
@@
which we can solve exactly via the Magnus expansion [PhysRevD.109.065024] to give
@@
After picking a probe state $\psi(q)$ on the basis of the pointer variable $Q$, using Eq (43) we can compute the Kraus operators for the field [PhysRevD.109.065024]
@@
In [ozawa1993canonical], Ozawa studies the process of taking such a limit, and shows that a measurement is only repeatable in that limit, in the sense that the outcomes of repeated measurements agree. Operationally, they consider allowing the interaction to continue for a longer time, say a factor of $T$, and rescaling the pointer units by $1/T$. It is shown in the case considered in that paper the standard deviation is scaled by $1/T$. This is mathematically the same as scaling the probe $x$ state’s argument by $T$, thus scaling $\sigma$ by $1/T$. We cannot do quite the same here, as we now discuss.
@@ IV.2.1 Regions and sharpness
To gain some intuition for the bound (50) on how accurately the field values can be probed locally, we can appeal to the results of [PhysRevD.109.065024], where it is shown that for $f$ given by a Gaussian with spacial width $\sigma_{S}$ and time width $\sigma_{T}$,
@@
In the $m\gg\frac{1}{\sigma_{T}}$ case, increasing either size of $\sigma_{T},\sigma_{S}$ makes the measurement sharper, while for $m\sim 0$, when $\sigma_{S}^{2}+\sigma_{T}^{2}$ is held constant, the region must be taken to be much wider spacially than temporally in order to get a sharp measurement. We see that concretely modelling a measurement leads to a channel that would not be easily guessed if avoiding signalling was the only concern. Indeed, channels comparable to ours have been guessed, as we will see in the next section, however they generally lack the bound on the sharpness we have found, which comes from the $\Delta_{R}(f,f)$ phase in the $S$-matrix. This phase arose, not from causality constraints, but from explicitly solving for the evolution of the field and pointer. As of such, it goes beyond Sorkin’s impossible measurements, as the channel is still causal without the $\Delta_{R}(f,f)$ phase. Interestingly, while the appearance of $\Delta_{R}(f,f)$ is not necessary for ruling out signalling of the field channel, it is key in the proof of continuous additivity and thus plays a role in the causality of the dilated channel, see appendix VII.3.
@@ IV.2.2 The model as a dilation
Finally, we comment on the connection to previous related work on families of Kraus operators and POVMs for field measurements. Our analysis provides a concrete dilation of the $L^{2}$-maps that defines families of Kraus operators in [albertini2023ideal], defined as
@@
This means that the ‘shape’ of the Kraus operators is specified by the initial state preparation $\psi$ of the probe. Similarly, there is a connection to the Kraus operators that are defined in [oeckl2024spectraldecompositionfieldoperators] in the case of instantaneous field measurements, that we will comment on in the next section. Lastly, in [mandrysch2024quantumfieldmeasurementsfewsterverch] the authors analyse a linear coupling between two equal mass real scalar field theories, building on [fewster2020quantum, Fewster2023]. By picking a squeezed quasi-local state for the probe field, they show that a Gaussian measurement of the probe leads to an instrument that in the weak coupling limit can be decomposed into a Gaussian measurement and a dephasing channel with certain strength. A full comparison between these results would require letting the masses of both fields vary independently. his is because our pointer is a single abstract degree of freedom, and has trivial internal dynamics, which would appear in a well controlled large mass limit of the probe represented as a quantum field, likely requiring us to reintroduce a probe Hamiltonian. We leave further exploration for future works.
@@ V Causality conditions on the Kraus operators
In this section we will examine the relevance of the continuous additivity Eq (31) for the concrete model of field measurements that we introduced in the previous section. We will demonstrate the model’s consistency with the framework laid out in section III, and derive the no-signalling and compositional properties of its Kraus operators, generalising those found in [oeckl2024spectraldecompositionfieldoperators]. These no-signalling identities ensure that this model does not induce impossible operations on the field.
@@
Identity (57) was proved in [oeckl2024spectraldecompositionfieldoperators] for an $\epsilon$-family of POVMs, including the spectral decomposition of $\phi(f)$. in the case of instantaneous operations defined over a Cauchy surface ($f$ being a purely spacial smearing) and it was proved to be sufficient for enforcing no-signalling condition Eq (21) in this case. Note that our results reduce to those in [oeckl2024spectraldecompositionfieldoperators] when we take $\psi$ to be Gaussian with width $\varepsilon$ and $f$ is supported on a spacial hypersurface so that $\Delta_{\textsc{r}}(f,f)=0$.
In our case, our concrete model describes time-extended operations, so (57) is not going to be sufficient to prove no-signalling and we will need the following additional identity [^16] The connection between this work and a generalised, i.e. ‘time extended’, version of Oeckl’s results in [oeckl2024spectraldecompositionfieldoperators] using a covariant path-integral description [oeckl2025causalmeasurementquantumfield] is work in progress, see also [oeckl2025localcompositionalmeasurementsquantum]..

### v2

@@ III.1.1 Eliminating the ‘impossible’
which is a coupling between the field and one probe, localised to some region $R_{i}$. Approaches to measurement theory in QFT in which the measurement process is modelled as a local scattering process and operations are induced by probes, include the detector-based measurement theory [39] and the Fewster-Verch (FV) framework in algebraic QFT [16]. In the detector-based approach, one works with (a general class) of Hamiltonian densities that model the detector-field coupling e.g. mimicking the light-matter interaction, and the local $S$-matrix can be defined as the time-ordered exponential of these interaction Hamiltonians (typically the predictions of the models rely on perturbation theory, using the Dyson expansion of the $S$-matrix). In the algebraic approach, the formulation of the FV framework relies on the structure of abstract *-algebras. In this setting, the interaction between system and probe field is formulated as a morphism $\Theta$ of the uncoupled theory, and it is not necessarily implemented through the adjoint action of a unitary scattering map $S$, namely as $\Theta(\cdot)=S^{\dagger}(\cdot)S$. Note that the approach we are taking here, is more abstract than the first (since it does not rely on particular interaction Hamiltonians and the Dyson expansion), but more concrete than the latter, since we assume the existence of these local scattering maps that implement the scatteting morphism in a Hilbert space.
@@
Partial or exact resolution to the impossible measurements scenario have been offered respectively in the detector-based approach [2, 14] and in the FV-framework [7]. In either case, an important ingredient of the resolution was an analogue causal factorisation property (32). An abstract version of causal factorisation is assumed to hold in the FV framework  [^11] The abstract version of the causal factorisation condition is that $\Theta_{\textsc{a+b}}=\Theta_{\textsc{b}}\circ\Theta_{\textsc{a}}$, which has to be proven for concrete models in this framework., while in [14] it was argued that that this property holds for a general class of Unruh-DeWitt-type detector models coupled locally to a coupled field over disjoint, causally orderable regions. Finally, all the measurement schemes considered in QFT correspond to so-called pre-measurement processes, involving only the unitary interaction between the field and the probe. Strictly speaking, the measurement is completed only once an outcome has been obtained, which would require, in turn, a measurement of the probe and so on, leading to an infinite regress, see [30]. In this work we ignore this final measurement stage and, following common jargon, refer to the pre-measurement stage itself as the “measurement process”.
@@
We should emphasise that in both approaches causal factorisation is assumed (or proven on a case-by-case basis) but, crucially, it is not sufficient for eliminating impossible measurements. More assumptions about the probe-field interactions in the coupling regions have to be made, i.e. the locality of the detector’s local current in the detector model approach [14]  [^12] In [14] the authors show that (seemingly local) Hamiltonian densities of the form $h(\mathsf{x})=\Lambda(\mathsf{x})J(\mathsf{x})\otimes\phi(\mathsf{x})$, where $J(\mathsf{x})$ is the current associated to a probe, can give rise to impossible measurements if the probe’s dynamical current is not microcausal (i.e. if $J(\mathsf{x}),J(\mathsf{y})]\neq 0$ for $\mathsf{x},\mathsf{y}$ spacelike separated) by using the Dyson expansion of $S=\mathcal{T}\text{exp}\,i\int h(\mathsf{x})\mathrm{d}\mathsf{x}$. , or the time-slice property of the coupled theory in the FV framework  [^13] The exact relation between continuous additivity and other dynamical axioms e.g. the time-slice property in the formal setting of (p)AQFT goes way beyond our purposes (for related results see e.g. [43]). [7]. Our claim is that Eq (31) captures the properties of a causal probe-field coupling that arises from the additional assumptions needed in the concrete and abstract frameworks, while making it clear that this can be directly related to internal factorisation properties of the interactions. Specifically, we will see that the ability to appropriately factorise the interaction region into causally ordered subregions is sufficient for ruling out impossible, i.e. signalling, operations. In the next section, we will see that it is not necessary, since a weaker factorisation property can be used.
@@ III.2 Hierarchy of factorisation conditions
We have introduced two notions of factorisations of $S$-matrices, causal factorisation and continuous additivity, that respectively correspond to composing disjoint interactions or decomposing a single interaction. Overall, there is a hierarchy of spacetime factorisation properties, as explained in [8]. In this section, we will briefly outline them here for completeness, and generalise our proof of no impossible operations to a weaker condition.
@@
One such weaker condition, called the Hammerstein property, is typically assumed in causal perturbation theory literature [42], where it is taken as continuous additivity Eq (31) may be too strong for very singular interactions [8]. The Hammerstein property is that
@@
Importantly, the Hammerstein property is also sufficient to rule out Sorkin problems, as we explicitly show in detail in Appendix VII.1 and sketch here.
@@
Looking at Figure 5, we can see a similar scenario to the proof for the stronger factorisation condition. As there, the supports of $f_{\pm}$ are spacelike to C,A respectively, and now $f_{0}$ is spacelike to A and C. This implies that $S[f_{+}+f_{0}]S[f_{0}]^{\dagger}$ commutes with $C$ and $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is localised spacelike to A. Following a very similar derivation to Eq (III.1.1) demonstrates that there is no signalling in a Sorkin scenario where $\Phi_{\textsc{b}}$ is induced by an $S$-matrix satisfying the Hammerstein property  [^14] Resolutions of the Sorkin problem in the abstract algebraic setting of the FV framework rely on the time-slice property of the coupled theory [7, 38]. At the level of the local $S$-matrix it has been shown that the time-slice property can be derived only from (36) and normalisation $S(0)=\openone [1$3], which explains why we were able to demonstrate the link between (36) and no impossible operations in Appendix VII.1.. For the full details, see Appendix VII.1.
@@ IV Operations induced by probes
In quantum mechanics, every unital, i.e. maps the identity to itself, CP map (UCP map) – that is, every non-selective operation on a quantum system – can be understood as arising from a unitary interaction between the system and an ancillary system, or probe, followed by discarding (tracing out) the probe. This is a consequence of the Stinespring dilation theorem, which ensures that any UCP transformation of the system can be “purified” by embedding it into a larger closed system that evolves unitarily. In the same vein, we will consider in this section the realisation of non-selective operations in QFT through interactions between the system (i.e., the field) and a probe. Then, a way to argue that unitary “kicks” which violate no-signalling are not physically realisable is to show that they cannot be implemented by any valid local interaction between the system and a probe (see also [31]).
@@ IV.1 Induced Kraus operators in QFT
The attempts of modelling local operations in QFT required a departure from the asymptotic scattering paradigm. In a series of papers [25, 26] Hellwig and Kraus model relativistic quantum measurement as a local scattering process. Their motivation for introducing the so-called Kraus operators, as induced through the dynamical coupling to some probe, was to explicate the operational interpretation of the local algebras that was implicit in the axiomatic approaches to QFT [23]. Their measurement scheme can be viewed as predecessor of recently proposed measurement frameworks in QFT [39, 16] that involve some (abstract or concrete) representation of dynamics between system and probe (for historical context see [19]).
@@
Kraus operators in scattering-based frameworks for local measurement in QFT have been derived previously in [46, 39, 37] for concrete Unruh-DeWitt-type models, and more recently in [30] for a concrete representation of the FV framework in a Hilbert space. Nevertheless, the operational consequences of the factorisation conditions at the level of Kraus operators have not been investigated. Here we investigate how these properties are captured by valid local scattering matrices and their (de)composition properties. Further, for certain concrete model for measurement of field amplitudes, we link these properties to a no-signalling property for the resulting Kraus operators.
@@ IV.2 A concrete model for local field measurements
We will use a simple model for field measurements (previously considered in [37]) where a scalar quantum field $\phi$ in Minkowski spacetime is coupled to a free pointer via $P$, with $[Q,P]=i$. We specify the following interaction
@@
Since we can pick any free Hamiltonian for the pointer that commutes with the interaction without much computational difficulty, we could in principle pick that of a relativistic scalar particle. However, to avoid any conceptual problems arising from interpreting the pointer as a relativistic particle—given the well known problems of causality and localisation in relativistic QM (e.g. Hegerfeldt’s [24] and Malament’s [29] theorems)—we elect simply to set $H_{\text{pointer}}=0$, viewing the pointer as an abstract, degree of freedom, not tied to spacetime. This has no effect on the field Kraus operators, and the change to the $S$-matrix can be found easily. To emphasise this conceptual choice, we will continue to denote spacetime points by $\mathsf{x}$ etc, while the internal values of the pointer as $q$. The interaction in (44) leads to the following scattering matrix
@@
which we can solve exactly via the Magnus expansion [37] to give
@@
After picking a probe state $\psi(q)$ on the basis of the pointer variable $Q$, using Eq (43) we can compute the Kraus operators for the field [37]
@@
In [36], Ozawa studies the process of taking such a limit, and shows that a measurement is only repeatable in that limit, in the sense that the outcomes of repeated measurements agree. Operationally, they consider allowing the interaction to continue for a longer time, say a factor of $T$, and rescaling the pointer units by $1/T$. It is shown in the case considered in that paper the standard deviation is scaled by $1/T$. This is mathematically the same as scaling the probe $x$ state’s argument by $T$, thus scaling $\sigma$ by $1/T$. We cannot do quite the same here, as we now discuss.
@@ IV.2.1 Regions and sharpness
To gain some intuition for the bound (50) on how accurately the field values can be probed locally, we can appeal to the results of [37], where it is shown that for $f$ given by a Gaussian with spacial width $\sigma_{S}$ and time width $\sigma_{T}$,
@@
In the $m\gg\frac{1}{\sigma_{T}}$ case, increasing either size of $\sigma_{T},\sigma_{S}$ makes the measurement sharper, while for $m\sim 0$, when $\sigma_{S}^{2}+\sigma_{T}^{2}$ is held constant, the region must be taken to be much wider spacially than temporally in order to get a sharp measurement. We see that concretely modelling a measurement leads to a channel that would not be easily guessed if avoiding signalling was the only concern. Indeed, channels comparable to ours have been guessed, as we will see in the next section, however they generally lack the bound on the sharpness we have found, which comes from the $\Delta_{\textsc{r}}(f,f)$ phase in the $S$-matrix. This phase arose, not from causality constraints, but from explicitly solving for the evolution of the field and pointer. As of such, it goes beyond Sorkin’s impossible measurements, as the channel is still causal without the $\Delta_{\textsc{r}}(f,f)$ phase. Interestingly, while the appearance of $\Delta_{\textsc{r}}(f,f)$ is not necessary for ruling out signalling of the field channel, it is key in the proof of continuous additivity and thus plays a role in the causality of the dilated channel, see Appendix VII.3.
@@ IV.2.2 The model as a dilation
Finally, we comment on the connection to previous related work on families of Kraus operators and POVMs for field measurements. Our analysis provides a concrete dilation of the $L^{2}$-maps that defines families of Kraus operators in [1], defined as
@@
This means that the ‘shape’ of the Kraus operators is specified by the initial state preparation $\psi$ of the probe. Similarly, there is a connection to the Kraus operators that are defined in [35] in the case of instantaneous field measurements, that we will comment on in the next section. Lastly, in [30] the authors analyse a linear coupling between two equal mass real scalar field theories, building on [16, 17]. By picking a squeezed quasi-local state for the probe field, they show that a Gaussian measurement of the probe leads to an instrument that in the weak coupling limit can be decomposed into a Gaussian measurement and a dephasing channel with certain strength. A full comparison between these results would require letting the masses of both fields vary independently. his is because our pointer is a single abstract degree of freedom, and has trivial internal dynamics, which would appear in a well controlled large mass limit of the probe represented as a quantum field, likely requiring us to reintroduce a probe Hamiltonian. We leave further exploration for future works.
@@ V Causality conditions on the Kraus operators
In this section we will examine the relevance of the continuous additivity Eq (31) for the concrete model of field measurements that we introduced in the previous section. We will demonstrate the model’s consistency with the framework laid out in section III, and derive the no-signalling and compositional properties of its Kraus operators, generalising those found in [35]. These no-signalling identities ensure that this model does not induce impossible operations on the field.
@@
Identity (57) was proved in [35] for an $\epsilon$-family of POVMs, including the spectral decomposition of $\phi(f)$. in the case of instantaneous operations defined over a Cauchy surface ($f$ being a purely spacial smearing) and it was proved to be sufficient for enforcing no-signalling condition Eq (21) in this case. Note that our results reduce to those in [35] when we take $\psi$ to be Gaussian with width $\varepsilon$ and $f$ is supported on a spacial hypersurface so that $\Delta_{\textsc{r}}(f,f)=0$.
In our case, our concrete model describes time-extended operations, so (57) is not going to be sufficient to prove no-signalling and we will need the following additional identity [^17] The connection between this work and a generalised, i.e. ‘time extended’, version of Oeckl’s results in [35] using a covariant path-integral description [34] is work in progress (see also [32])..